# Flavor diagonal nucleon charges using clover fermions on MILC HISQ ensembles

### v2

@@ Appendix A Chiral-continuum-finite-volume (CCFV) extrapolation for $g_{A,S,T}^{u,d,s}$
Figures 816 in this appendix show the CCFV fits used to extrapolate the data for the charges on the eight ensembles to the physical point. In each case, the three panels show the simultaneous CCFV fit (pink band) plotted versus either $a$ or $M_{\pi}^{2}$ or $M_{\pi}L$, respectively, with the other two variables set to their physical point values. In the panels plotted versus $M_{\pi}^{2}$, the additional grey bands show the fit with only the chiral corrections included in Eqs. (7) or (8). Note that when the discretization and finite volume corrections are small, the grey bands overlap with the CCFV fits (pink bands). The results of these fits are summarized in Table 6.
Figure 8: Chiral-continuum-finite-volume (CCFV) extrapolation for $g_{A}^{u}$ using Eq. (7). The 3 panels in each column show the result of the simultaneous fit (pink bands) plotted versus $a$, $M_{\pi}^{2}$ and $M_{\pi}L$, respectively, with the other two variables set to their physical point values. The grey band shows a chiral fit, i.e., assuming $c_{1}=c_{3}=0$ in Eq. (7). Each row shows, from left to right, the fits for the four different analysis: (i) {standard,$\rm{Z_{1}}$}, (ii) {standard,$\rm{Z_{2}}$}, (iii) $N\pi$,$\rm{Z_{1}}$} and (iv) $N\pi$,$\rm{Z_{2}}$}. The red star is the value at the physical point given by the CCFV fit and the black cross is the result of just the chiral fit.
@@
Figure 13: CCFV extrapolation for $g_{T}^{s}$ using Eq. (7). The rest is the same as in Fig. 8. These data are only analyzed using the “standard” strategy as the likely multihadron excited state, $\Sigma K$, is heavier than the $N(1440)$ radial excitation.
@@
Figure 17: The Chiral-continuum (CC) extrapolation for $\sigma_{\pi N}=m_{ud}g_{S}^{u+d}$ using the ansatz given in Eq. (9). The label $3^{\chi}$ implies that the coefficient $d_{3}^{\chi}$ is set to the $\chi$PT value taken from Ref. [4] The other four parameters are outputs of the fit. The color (grey/black) bands and points correspond to the “$N\pi$” (“standard”) analysis.
@@ Appendix B Nonperturbative renormalization for the 2+1-flavor theory using the RI-sMOM scheme
This Appendix presents an overview of the calculation of the renormalization constants for the flavor diagonal quark bilinear operators $\mathcal{O}^{f}=\bar{\psi}^{f}\Gamma\psi^{f}$ with Dirac matrix $\Gamma$ and the flavor index $f=\{u,d,s\}$ in the $N_{f}=3$ theory. The lattice calculation is done in the RI-sMOM scheme [8, 9] on four ensembles at different values of the lattice spacing used in this study: $a15m310$, $a12m310$, $a09m310$, and $a06m310$. Since data are obtained at only one value of $M_{\pi}\neq 0$ at each $a$, we explicitly assume that the quark mass dependence can be neglected and these results are a good approximation to the chiral limit values. Second, the lattice calculation is not fully $O(a)$ improved in either the action or the operators. In the latter, we neglect all the lattice spacing dependent improvement terms. Consequently, the results for the renormalized charges are extrapolated to the continuum limit using an ansatz starting with a term linear in $a$.
Results for the matching factor between the RI-sMOM and the ${\overline{\text{MS}}}$ scheme, and the running in the ${\overline{\text{MS}}}$ scheme are taken from continuum perturbation theory [62, 63, 64, 65, 66, 67] as specified in Appendix B.3.
The general relation between renormalized, $\mathcal{O}_{R}$, and bare, $\mathcal{O}$, operators including mixing between flavors is given by
$\mathcal{O}^{f}_{R}$ $=\sum_{f^{\prime}}Z_{\Gamma}^{ff^{\prime}}\mathcal{O}^{f^{\prime}}\,.$ (11)
We determine $Z_{\Gamma}^{ff^{\prime}}$ nonperturbatively on the lattice using the regularization independent (RI) renormalization scheme [8] in which the renormalized vertex function is set to its tree-level value. The calculation is done with the gauge fields fixed to the Landau gauge and quark propagators generated using momentum sources. The final results for the three charges, using the two methods, labeled $\rm{Z_{1}}$ and $\rm{Z_{2}}$ and discussed in Appendix B.4, for partially removing discretization errors are given in Appendix B.5.
@@ B.1 Flavor mixing in the RI scheme
The (-) sign in $\Gamma_{\text{disc}}$ is due to the anticommuting nature of the fermion fields, i.e., for the quark loop. On the lattice we calculate the projected amputated vertex functions $c_{f}$ (connected) and $d_{ff^{\prime}}$ (disconnected) defined as
@@
For the discussion of the matching and running, the more convenient basis is the isovector (3), octet (8), and singlet (0) one, i.e., $f\in\{u-d,u+d-2s,u+d+s\}$. In that case the renormalization factors are
$\left(\begin{array}[]{ccc}Z^{3,3}_{\Gamma}&0&0\\ 0&Z^{8,8}_{\Gamma}&Z^{8,0}_{\Gamma}\\ 0&Z^{0,8}_{\Gamma}&Z^{0,0}_{\Gamma}\end{array}\right)=\left(\begin{array}[]{ccc}c_{\Gamma}^{l}&0&0\\ 0&\frac{1}{3}(c_{\Gamma}^{l}+2(c_{\Gamma}^{s}-d_{\Gamma}^{ll}+d_{\Gamma}^{ls}+d_{\Gamma}^{sl}-d_{\Gamma}^{ss}))&\frac{2}{3}(c_{\Gamma}^{l}-c_{\Gamma}^{s}-2d_{\Gamma}^{ll}+2d_{\Gamma}^{ls}-d_{\Gamma}^{sl}+d_{\Gamma}^{ss})\\ 0&\frac{1}{3}(c_{\Gamma}^{l}-c_{\Gamma}^{s}-2d_{\Gamma}^{ll}-d_{\Gamma}^{ls}+2d_{\Gamma}^{sl}+d_{\Gamma}^{ss})&\frac{1}{3}(2c_{\Gamma}^{l}+c_{\Gamma}^{s}-4d_{\Gamma}^{ll}-2d_{\Gamma}^{ls}-2d_{\Gamma}^{sl}-d_{\Gamma}^{ss})\end{array}\right)^{-1}$ (27)
@@ B.3 Matching and RG running
$Z^{match}_{S,\text{SMOM}}$ $=$ $1+0.64552a\ +\ (23.024\,-4.0135n_{f})a^{2}$
@@
$Z^{match}_{T,\text{SMOM}}$ $=$ $1-0.215a\ +\ (4.103n_{f}-43.384)a^{2}$ (31)
@@
$C_{A}^{\text{RI-sMOM}\to{\overline{\text{MS}}}}(\mu)=\left(\begin{array}[]{ccc}1&0&0\\ 0&1&0\\ 0&0&1-26.5\,a^{2}\,n_{f}\end{array}\right)$ (35)
@@
On the lattice, Wilson-clover fermions break chiral symmetry explicitly and there are non-zero contributions from disconnected loops for all charges. These give rise to off-diagonal terms as shown in Eqs. (20) and (27). The mixing in the tensor channel is $O(\alpha_{s}^{3})$ in lattice perturbation theory [74] implying $r_{T}\equiv Z_{T}^{\text{singlet}}/Z_{T}^{\text{non-singlet}}=1+O(\alpha_{s}^{3})\approx 1$ where $a=\alpha_{s}/4\pi$ and $n_{f}$ denotes the number of flavors. The largest disconnected contribution (off-diagonal elements $d_{\Gamma}^{ff^{\prime}}$) is for the scalar. It is an approximately $4\%$ effect for the coarsest $a\approx 0.15$ fm ensemble and below 1% on the rest.
@@ B.4 Two renormalization strategies $\rm{Z_{1}}$ and $\rm{Z_{2}}$ defined by the choice of $Z_{\psi}^{q}$
These two strategies were used in Refs. [70, 19] for isovector bilinear operators of light quarks, i.e., $g_{\Gamma}Z_{\Gamma}$ and $Z_{\Gamma}/Z_{V}\times g_{\Gamma}/g_{V}$. Later, they were called $\rm Z_{1}$ and $\rm Z_{2}$ in Ref. [11]. In Refs. [70, 19] we showed that they have different behavior versus $q^{2}$ as the discretization effects are different in $Z_{\psi}$ and $Z_{\psi}^{\rm VWI}$, which are obtained from $S_{B}(p)$ and $c_{\Gamma}^{V}(p)$, respectively333 The $O(a)$ improvement for $S_{B}(p)$ is studied in Ref. [75], and $c_{\Gamma}^{V}(p)$ (off-shell improvement) is studied in Ref. [76]. To within expected discretization errors, we checked

### v3

@@ Appendix A Chiral-continuum-finite-volume (CCFV) extrapolation for $g_{A,S,T}^{u,d,s}$
Figures 816 in this appendix show the CCFV fits used to extrapolate the data for the charges on the eight ensembles to the physical point. In each column, the three panels show the simultaneous CCFV fit (pink band) plotted versus either $a$ or $M_{\pi}^{2}$ or $M_{\pi}L$, respectively, with the other two variables set to their physical point values. In the panels plotted versus $M_{\pi}^{2}$, the additional grey bands show the fit with only the chiral corrections included in Eqs. (7) or (8). Note that when the discretization and finite volume corrections are small, the grey bands overlap with the CCFV fits (pink bands). The results of these fits are summarized in Table 6.
[…]
@@
Figure 13: CCFV extrapolation for $g_{T}^{s}$ using Eq. (7). The rest is the same as in Fig. 8. These data are only analyzed using the “standard” strategy as the likely multihadron lowest excited state, $\Sigma K$, is heavier than the $N(1440)$ radial excitation.
@@
[…]
@@ Appendix B Nonperturbative renormalization for the 2+1-flavor theory using the RI-sMOM scheme
This Appendix presents an overview of the calculation of the renormalization constants for the flavor diagonal quark bilinear operators $\mathcal{O}^{f}=\bar{\psi}^{f}\Gamma\psi^{f}$ with Dirac matrix $\Gamma$ and the flavor index $f=\{u,d,s\}$ in the $N_{f}=3$ theory. The lattice calculation is done in the RI-sMOM scheme [8, 9] on four ensembles at the different values of the lattice spacing used in this study: $a15m310$, $a12m310$, $a09m310$, and $a06m310$. Since data are obtained at only one value of $M_{\pi}\neq 0$ at each $a$, we explicitly assume that the quark mass dependence can be neglected and these results are a good approximation to the chiral limit values. Second, the lattice calculation is not fully $O(a)$ improved in neither the action nor the operators. In the latter, we neglect all the lattice spacing dependent improvement terms. Consequently, the results for the renormalized charges are extrapolated to the continuum limit using an ansatz starting with a term linear in $a$.
Results for the matching factors between the RI-sMOM and the $\overline{\text{MS}}$} scheme, and the running in the $\overline{\text{MS}}$} scheme are taken from continuum perturbation theory [62, 63, 64, 65, 66, 67] as specified in Appendix B.3.
The general relation between renormalized, $\mathcal{O}_{R}$, and bare, $\mathcal{O}$, operators including mixing between flavors is:
$\mathcal{O}^{f}_{R}$ $=\sum_{f^{\prime}}Z_{\Gamma}^{ff^{\prime}}\mathcal{O}^{f^{\prime}}\,$, (11)
with the matrix $Z_{\Gamma}^{ff^{\prime}}$ determined nonperturbatively on the lattice using the regularization independent (RI) renormalization scheme [8] in which the renormalized vertex function is set to its tree-level value. The calculation is done with the gauge fields fixed to the Landau gauge and quark propagators generated using momentum sources. The final results for the three charges, using the two methods, labeled $\rm{Z_{1}}$ and $\rm{Z_{2}}$ and discussed in Appendix B.4, for partially removing discretization errors are given in Appendix B.5.
@@ B.1 Flavor mixing in the RI scheme
The (-) sign in $\Gamma_{\text{disc}}$ is due to the anticommuting nature of the fermion fields contracted into the quark loop. On the lattice we calculate the projected amputated vertex functions $c_{f}$ (connected) and $d_{ff^{\prime}}$ (disconnected) defined as
@@
For the discussion of the matching and running, the more convenient basis is the isovector (3), octet (8), and singlet (0) one, i.e., $f\in\{u-d,u+d-2s,u+d+s\}$. In this case the renormalization matrix in terms of the 6 quantities is
$\hskip-22.76228pt\left(\begin{array}[]{ccc}Z^{3,3}_{\Gamma}&0&0\\ 0&Z^{8,8}_{\Gamma}&Z^{8,0}_{\Gamma}\\ 0&Z^{0,8}_{\Gamma}&Z^{0,0}_{\Gamma}\end{array}\right)=\left(\begin{array}[]{ccc}c_{\Gamma}^{l}&0&0\\ 0&\frac{1}{3}(c_{\Gamma}^{l}+2(c_{\Gamma}^{s}-d_{\Gamma}^{ll}+d_{\Gamma}^{ls}+d_{\Gamma}^{sl}-d_{\Gamma}^{ss}))&\frac{2}{3}(c_{\Gamma}^{l}-c_{\Gamma}^{s}-2d_{\Gamma}^{ll}+2d_{\Gamma}^{ls}-d_{\Gamma}^{sl}+d_{\Gamma}^{ss})\\ 0&\frac{1}{3}(c_{\Gamma}^{l}-c_{\Gamma}^{s}-2d_{\Gamma}^{ll}-d_{\Gamma}^{ls}+2d_{\Gamma}^{sl}+d_{\Gamma}^{ss})&\frac{1}{3}(2c_{\Gamma}^{l}+c_{\Gamma}^{s}-4d_{\Gamma}^{ll}-2d_{\Gamma}^{ls}-2d_{\Gamma}^{sl}-d_{\Gamma}^{ss})\end{array}\right)^{-1}\,.$ (27)
@@ B.3 Matching and RG running
$\hskip-17.07182ptZ^{match}_{S,\text{SMOM}}$ $=$ $1+0.64552a\ +\ (23.024\,-4.0135n_{f})a^{2}$
@@
$\hskip-17.07182ptZ^{match}_{T,\text{SMOM}}$ $=$ $1-0.215a\ +\ (4.103n_{f}-43.384)a^{2}$ (31)
@@
$\hskip-19.91684ptC_{A}^{\text{RI-sMOM}\to{\overline{\text{MS}}}}(\mu)=\left(\begin{array}[]{ccc}1&0&0\\ 0&1&0\\ 0&0&1-26.5\,a^{2}\,n_{f}\end{array}\right)$ (35)
@@
On the lattice, Wilson-clover fermions break chiral symmetry explicitly and there are non-zero contributions from disconnected loops for all charges. These give rise to off-diagonal terms as shown in Eqs. (20) and (27). The mixing in the tensor channel is $O(\alpha_{s}^{3})$ in lattice perturbation theory [74] implying $r_{T}\equiv Z_{T}^{\text{singlet}}/Z_{T}^{\text{non-singlet}}=1+O(\alpha_{s}^{3})\approx 1$ where $a=\alpha_{s}/4\pi$ and $n_{f}$ denotes the number of flavors. The largest disconnected contribution (off-diagonal elements $d_{\Gamma}^{ff^{\prime}}$) is for the axial. In the $\{u-d,u+d,s\}$ basis, it is approximately $7\%$ for the coarsest $a\approx 0.15$ fm ensemble and decreases to below 2% at the finest $a\approx 0.06$ fm ensemble.
@@ B.4 Two renormalization strategies $\rm{Z_{1}}$ and $\rm{Z_{2}}$ defined by the choice of $Z_{\psi}^{q}$
These two strategies were used in Refs. [70, 19] for isovector bilinear operators of light quarks, i.e., $g_{\Gamma}Z_{\Gamma}$ and $Z_{\Gamma}/Z_{V}\times g_{\Gamma}/g_{V}$. Later, they were called $\rm Z_{1}$ and $\rm Z_{2}$ in Ref. [11]. In Refs. [70, 19] we showed that they have different behavior versus $q^{2}$ as the discretization effects are different in $Z_{\psi}$ and $Z_{\psi}^{\rm VWI}$, which are obtained from $S_{B}(p)$ and $c_{\Gamma}^{V}(p)$, respectively [^3] The $O(a)$ improvement for $S_{B}(p)$ is studied in Ref. [75], and $c_{\Gamma}^{V}(p)$ (off-shell improvement) is studied in Ref. [76]. To within expected discretization errors, we checked